# DiFlowDubber: Discrete Flow Matching for Automated Video Dubbing via Cross-Modal Alignment and Synchronization

### 17 Mar 2026 v2

@@ Abstract
Video dubbing has broad applications in filmmaking, multimedia creation, and assistive speech technology. Existing approaches either train directly on limited dubbing datasets or adopt a two-stage pipeline that adapts pre-trained text-to-speech (TTS) models, which often struggle to produce expressive prosody, rich acoustic characteristics, and precise synchronization. To address these issues, we propose DiFlowDubber with a novel two-stage training framework that effectively transfers knowledge from a pre-trained TTS model to video-driven dubbing, with a discrete flow matching generative backbone. Specifically, we design a FaPro module that captures global prosody and stylistic cues from facial expressions and leverages this information to guide the modeling of subsequent speech attributes. To ensure precise speech-lip synchronization, we introduce a Synchronizer module that bridges the modality gap among text, video, and speech, thereby improving cross-modal alignment and generating speech that is temporally synchronized with lip movements. Experiments on two primary benchmark datasets demonstrate that DiFlowDubber outperforms previous methods across multiple metrics. The demo page is available at: https://nngocson2002.github.io/projects/diflowdubber.
@@ 1 Introduction
Figure 1: Overall inference pipeline of DiFlowDubber. The Face-to-Prosody Mapper module predicts global prosody priors that capture global prosody and stylistic cues from facial expressions. The Content-Consistent Temporal Adaptation module generates discrete content tokens conditioned on lip movements, text, and prosody priors, ensuring consistent with the target text transcription and temporal alignment. Discrete Flow-based Prosody-Acoustic module generates diverse yet globally consistent prosody tokens under the guidance of the prosody prior, together with corresponding acoustic tokens. The speech waveform is synthesized from the predicted tokens and speaker embedding via a Codec Decoder.
@@
In the first stage, we develop a zero-shot TTS system that models factorized speech representations to synthesize high-quality speech in a discrete setting. To this end, the system is pre-trained in a zero-shot TTS configuration, where we employ a simple deterministic content modeling architecture to efficiently capture linguistic structure. In contrast, for prosody and acoustic attributes, we adopt a generative modeling approach using discrete flow matching through the Discrete Flow-Based Prosody-Acoustic (DFPA) module to capture expressive prosodic variations and realistic acoustic diversity from the corpus. In the second stage, the pre-trained TTS model is adapted for video dubbing, as illustrated in Figure 1. To address the synchronization challenge during domain adaptation, we introduce the Synchronizer, which takes phoneme sequences and lips-cropped visual features as inputs to capture temporally aligned content details. In addition, we design the Face-to-Prosody Mapper (FaPro) module to extract global prosodic cues from face-cropped video frames. These features are fused with the synchronized content representation to construct a rich, fine-grained multimodal embedding that captures prosody-content correlations. Finally, this embedding, together with the global prosody features, is fed into the DFPA, which is initialized and fine-tuned from the pre-trained TTS model to generate diverse prosodic variations along with the corresponding acoustic tokens. We summarize our contributions as follows:
  • We propose DiFlowDubber, an end-to-end automated video dubbing framework built upon a DFM generative backbone. Our approach consists of a zero-shot TTS pre-training stage that fully exploits content information, expressive prosodic variations, and realistic acoustic diversity from large-scale corpora, followed by a transfer learning stage that adapts the pretrained TTS model to visual inputs for video dubbing. To the best of our knowledge, this is the first work to extend DFM to video dubbing.
  • We propose a novel Synchronizer module that effectively bridges the modality gap between text, video, and speech. This module enhances video-speech alignment through a dual-alignment mechanism which collectively enforces robust cross-modal synchronization.
@@ 2.1 Speech synthesis
Speech synthesis has undergone rapid progress in recent years, moving from concatenative and statistical parametric approaches [26, 61] to deep generative models that achieve human-like naturalness. Early neural architectures [58, 51] demonstrated the feasibility of end-to-end TTS by jointly modeling prosody and waveform generation. Subsequent advances have focused on improving efficiency and controllability, with non-autoregressive models [50, 31, 48] reducing inference latency while maintaining quality. More recently, discrete representation learning [1, 46, 29, 4, 57] and diffusion/flow-based methods [31, 29, 20, 6, 24, 41, 30, 52, 28, 34, 32, 33, 38, 17] have further improved expressiveness and robustness, allowing fine-grained control over prosody, speaker identity, and style. Parallel to these developments, large language models have been integrated with TTS pipelines [22, 18, 65, 21, 3, 53, 39, 44, 27, 56, 5], bridging the gap between linguistic understanding and acoustic realization. Despite these advances, most systems remain text-driven, limiting their applicability in multimodal scenarios such as video dubbing, where visual articulatory cues are essential for precise synchronization. Building on these foundations, we extend discrete generative modeling to incorporate visual cues, addressing the modality gap between text-based TTS and video-conditioned speech synthesis.
@@ 3 Methodology
Figure 2: Pipeline of the proposed DiFlowDubber. Our framework comprises a two-stage pipeline. The first stage performs zero-shot TTS pre-training, where a simple deterministic content modeling architecture efficiently captures linguistic structure (orange dashed box). For prosody and acoustic attributes, we adopt the (b) Discrete Flow-Based Prosody-Acoustic (DFPA) module to model expressive prosodic variations and realistic acoustic diversity from the corpus. In the second stage, the model is adapted to V2C task. The (a) Content-Consistent Temporal Adaptation (CCTA) module transfers consistent content knowledge from the TTS domain and generates temporally aligned content representations, while the FaPro module extracts a global prosody prior from facial expression cues. The DFPA module then models the joint distribution of prosody and acoustic tokens conditioned on the prosody prior and latent content representations.
@@ 3.1 Overview
Our framework is designed as an end-to-end pipeline. We first pre-train a zero-shot TTS system that learns to control and model complex speech attributes, including prosody, content, and acoustic characteristics. The content attribute is modeled as illustrated in Figure 2 (Teacher box), while the Discrete Flow-Based Prosody-Acoustic (DFPA) generates tokens that capture expressive prosody and realistic acoustic characteristics consistent with a given speech prompt (Figure 2b, orange arrow). We then adapt this pre-trained model to the movie dubbing domain, enabling it to inherit the advantages of the large-scale TTS corpus by integrating visual cues, as illustrated in Figure 1. This adaptation is achieved through two key components: (a) Content-Consistent Temporal Adaptation (CCTA) (Section 3.4.3), which transfers knowledge from the TTS domain to generate content representations consistent with the target text transcription, achieving precise lip synchronization and capturing captures global prosodic states conveyed through visual cues; and (b) DFPA modeling with modified conditioning to better suit the visual modality which is detailed in Section 3.4.4.
@@ 3.2 Data Preparation
We leverage the pre-trained FACodec [28], which consists of a Codec Encoder and a Codec Decoder. The encoder decomposes raw speech into prosody, content, and acoustic token sequences, which serve as our discrete target data, along with a speaker identity embedding. The decoder subsequently reconstructs the speech waveform from these tokens and the corresponding speaker identity.
Given a speech waveform $\mathbf{S}$, we apply the Codec Encoder to extract a sequence of discrete speech tokens along with a speaker embedding. This process yields $\mathbf{x}_{p}\in\mathbb{R}^{m\times L}$, $\mathbf{x}_{c}\in\mathbb{R}^{n\times L}$, $\mathbf{x}_{a}\in\mathbb{R}^{k\times L}$, and $\mathbf{s}\in\mathbb{R}^{D_{s}}$, representing the prosody, content, and acoustic tokens, and the speaker embedding, respectively. Here, $m$, $n$, and $k$ denote the number of residual vector quantization (RVQ) codebooks for prosody, content, and acoustic representations, respectively, each with a vocabulary size of $v$. $L$ is the token sequence length, and $D_{s}$ is the speaker embedding dimension.
@@
Given a text transcript $\mathbf{T}$, we use a grapheme-to-phoneme converter [42] to obtain the corresponding phoneme sequence, which is mapped into a sequence of tokens $\mathbf{P}=(\mathbf{P}_{1},\mathbf{P}_{2},\dots,\mathbf{P}_{N})$ consisting of $N$ tokens. A phoneme encoder then transforms $\mathbf{P}$ into a sequence of embeddings $\mathbf{p}=(\mathbf{p}_{1},\mathbf{p}_{2},\dots,\mathbf{p}_{N})\in\mathbb{R}^{N\times D}$, where $D$ denotes the hidden dimension.
@@
Given a silent video $\mathbf{V}$, we follow the extraction pipeline in [12] to obtain the face-cropped video $\mathbf{V}_{\text{face}}$ and lips-cropped video $\mathbf{V}_{\text{lip}}$ and their corresponding face feature embeddings $\mathbf{v}_{\text{face}}\in\mathbb{R}^{F\times D_{f}}$ and lip feature embeddings $\mathbf{v}_{\text{lip}}\in\mathbb{R}^{F\times D_{l}}$. Here, $F$ is the number of frames, and $D_{f}$ and $D_{l}$ denote the face and lip feature dimensions, respectively. Since these extracted features exist in different representation spaces, we apply a 1D convolution to map them into a shared space: $\mathbf{v}_{\text{face}}\in\mathbb{R}^{F\times D}$ and $\mathbf{v}_{\text{lip}}\in\mathbb{R}^{F\times D}$.
@@ 3.3 Zero-shot TTS Pre-training
This stage aims to generate high-quality speech with natural prosody and expressive characteristics. To achieve this, we follow the zero-shot TTS setting, which aims to synthesize speech that faithfully reproduces the speech attributes of unseen speakers using only a few seconds of their reference speech. Let $\mathbf{S}_{r}$ denote the reference speech. Our objective is to build a framework that synthesizes $\tilde{\mathbf{S}}$ by generating prosody $\tilde{\mathbf{x}}_{p}$, content $\tilde{\mathbf{x}}_{c}$, and acoustic $\tilde{\mathbf{x}}_{a}$ discrete token sequences conditioned on $\mathbf{S}_{r}$ and $\mathbf{T}$. The key components of this stage are described below.
@@
As shown in Figure 2, we model the content attribute using a simple deterministic architecture (Teacher box). The phoneme sequence $\mathbf{p}$ is processed to predict $n$ representations for each residual codebook of the content tokens. These representations are then used to produce a content latent representation $\tilde{\mathbf{h}}_{c}$ and a content probability distribution employed to predict the content tokens $\tilde{\mathbf{x}}_{c}$.
@@
For prosody and acoustic attributes, we design a discrete flow-matching framework to model their joint distribution. As illustrated in Figure 2b, the denoiser outputs the joint prosody-acoustic distribution, represented as a probabilistic velocity field within the discrete flow matching process. Our denoiser is built upon a Diffusion Transformer (DiT) [43], which takes the current denoising target $\mathbf{x}_{t}$, obtained by concatenating the masked prosody and acoustic token sequences, as input. It is conditioned on the content latent representation $\tilde{\mathbf{h}}_{c}$, prosody tokens, acoustic tokens, and a speaker embedding extracted from $\mathbf{S}_{r}$, serving as contextual conditioning to predict features in $\mathbb{R}^{(m+k)\times L\times v}$ that represent the posterior distributions for prosody and acoustic tokens. This design enables the model to simultaneously generate tokens that capture expressive prosody $\tilde{\mathbf{x}}_{p}$ and realistic acoustic characteristics $\tilde{\mathbf{x}}_{a}$, remaining consistent with the speaker’s vocal identity. The objective of the discrete flow matching process is to learn a probabilistic denoiser $p(\mathbf{x}_{1}|\mathbf{x}_{t})$ that reconstructs masked tokens under varying masking ratios. Formally, the model is optimized to minimize the cross-entropy loss:
@@
where $p_{1|t}(\cdot\mid\mathbf{x}_{t},\mathbf{c};\theta)$ denotes the posterior distribution of $\mathbf{x}_{1}$ given the partially corrupted sequence $\mathbf{x}_{t}$, conditioning inputs $\mathbf{c}$, and $\theta$ represents the parameters of the discrete flow matching prediction network. Here, $\mathcal{T}=\{1,\ldots,(m+k)L\}$ represents the index set of token positions, $\mathbf{x}_{t}\sim p_{t}(\mathbf{x}\mid\mathbf{x}_{0},\mathbf{x}_{1})$ is sampled from the mixture path, and $\mathbf{x}_{0}\sim p$, $\mathbf{x}_{1}\sim q$ denote the source and target distributions, respectively.
Once trained, this TTS model serves as a teacher in our dubbing framework. Its pre-learned representations of pronunciation, prosody, and acoustic consistency are transferred to the video-driven dubbing stage, providing strong initialization for naturalness, pronunciation accuracy, and expressive speech generation. We encourage readers refer to the Supplementary Material for detailed architecture and training configurations of this stage.

### 03 Apr 2026 v4

@@ Abstract
Video dubbing requires content accuracy, expressive prosody, high-quality acoustics, and precise lip synchronization, yet existing approaches struggle on all four fronts. To address these issues, we propose DiFlowDubber, the first video dubbing framework built upon a discrete flow matching backbone with a novel two-stage training strategy. In the first stage, a zero-shot text-to-speech (TTS) system is pre-trained on large-scale corpora, where a deterministic architecture captures linguistic structures, and the Discrete Flow-based Prosody-Acoustic (DFPA) module models expressive prosody and realistic acoustic characteristics. In the second stage, we propose the Content-Consistent Temporal Adaptation (CCTA) to transfer TTS knowledge to the dubbing domain: its Synchronizer enforces cross-modal alignment for lip-synchronized speech. Complementarily, the Face-to-Prosody Mapper (FaPro) conditions prosody on facial expressions, whose outputs are then fused with those of the Synchronizer to construct rich, fine-grained multimodal embeddings that capture prosody-content correlations, guiding the DFPA to generate expressive prosody and acoustic tokens for content-consistent speech. Experiments on two benchmark datasets demonstrate that DiFlowDubber outperforms prior methods across multiple evaluation metrics.
@@ 1 Introduction
Figure 1: Overall Inference Pipeline of DiFlowDubber. Face-to-Prosody Mapper module predicts global prosody priors that capture global prosody and stylistic cues from facial expressions. Content-Consistent Temporal Adaptation module generates discrete content tokens conditioned on lip movements, text, and prosody priors, ensuring consistency with the target text transcription and temporal alignment. Discrete Flow-based Prosody-Acoustic module generates diverse yet globally consistent prosody tokens under the guidance of the prosody prior, together with corresponding acoustic tokens. The speech waveform is synthesized from the predicted tokens and speaker embedding via a Codec Decoder.
@@
In the first stage, we develop a zero-shot TTS system that models factorized speech representations to synthesize high-quality speech in a discrete setting. To this end, the system is pre-trained in a zero-shot TTS configuration, where we employ a simple deterministic content modeling architecture to efficiently capture linguistic structure. In contrast, for prosody and acoustic attributes, we adopt a generative modeling approach using discrete flow matching through the Discrete Flow-based Prosody-Acoustic (DFPA) module to capture expressive prosodic variations and realistic acoustic diversity from the corpus. In the second stage, the pre-trained TTS model is adapted for video dubbing, as illustrated in Figure 1. In particular, we first design the Face-to-Prosody Mapper (FaPro) module to extract global prosodic cues from face-cropped video frames. In addition, we introduce the Content-Consistent Temporal Adaptation (CCTA) module, which transfers knowledge from the TTS domain to generate content representations consistent with the target text transcription. Central to this module is the Synchronizer, which replaces the duration predictor of the TTS model and takes phoneme sequences and lip-cropped visual features as inputs to capture temporally aligned content details, thereby addressing the synchronization challenge during domain adaptation. The global prosody prior from FaPro and the synchronized content representation from the Synchronizer module are then fused to construct rich, fine-grained multimodal embeddings that capture prosody-content correlations. Finally, this embedding is used to predict content token sequences and, together with the global prosody features, is fed into the DFPA, which is initialized from the pre-trained TTS model and fine-tuned to generate diverse prosody and acoustic tokens. We summarize our contributions as follows:
  • We propose DiFlowDubber, the first end-to-end automated video dubbing framework built upon a discrete flow matching generative backbone. Our approach consists of a zero-shot TTS pre-training stage that fully exploits content information, expressive prosodic variations, and realistic acoustic diversity from large-scale corpora, followed by a transfer learning stage that adapts the pretrained TTS model to visual inputs for video dubbing.
  • We propose the CCTA module for effective knowledge transfer from the TTS domain to the video dubbing setting. Central to CCTA is the Synchronizer, which replaces the duration predictor of the TTS model with a dual-alignment mechanism over text, video, and speech, enforcing precise cross-modal synchronization while preserving content consistency with the target transcription. Complementarily, the FaPro extracts global prosodic cues from facial expressions to guide expressive speech generation.
@@ 2.1 Speech synthesis
Speech synthesis has undergone rapid progress in recent years, moving from concatenative and statistical parametric approaches [25, 58] to deep generative models that achieve human-like naturalness. Early neural architectures [55, 48] demonstrated the feasibility of end-to-end TTS by jointly modeling prosody and waveform generation. Subsequent advances have focused on improving efficiency and controllability, with non-autoregressive models [47, 29, 46] reducing inference latency while maintaining quality. More recently, discrete representation learning [1, 44, 27, 4, 54] and diffusion/flow-based methods [29, 19, 5, 23, 39, 28, 49, 27, 32, 30, 31, 36, 16] have further improved expressiveness and robustness, allowing fine-grained control over prosody, speaker identity, and style. Parallel to these developments, large language models have been integrated with TTS pipelines [21, 17, 62, 20, 3, 50, 37, 42, 26, 53, 4], bridging the gap between linguistic understanding and acoustic realization. Despite these advances, most systems remain text-driven, limiting their applicability in multimodal scenarios such as video dubbing, where visual articulatory cues are essential for precise synchronization. Building on these foundations, we extend discrete generative modeling to incorporate visual cues, addressing the modality gap between text-based TTS and video-conditioned speech synthesis.
@@ 3 Methodology
Figure 2: Detailed of DiFlowDubber Pipeline. Our framework is a two-stage pipeline. The first stage performs zero-shot TTS pre-training, where a simple deterministic content modeling architecture efficiently captures linguistic structures (orange dashed box). For prosody and acoustic attributes, we adopt (b) Discrete Flow-based Prosody-Acoustic (DFPA) module to model expressive prosodic variations and realistic acoustic diversity from the corpus. In the second stage, the model is adapted to the video dubbing task, where (a) Content-Consistent Temporal Adaptation module transfers consistent content knowledge from the TTS domain and generates temporally aligned content representations. Meanwhile, FaPro extracts a global prosody prior from facial expression cues. The DFPA module then models the joint distribution of prosody and acoustic tokens conditioned on the global prosody prior, latent content representations, and speaker embedding.
@@ 3.1 Overview
Our framework is designed as an end-to-end pipeline. We first pre-train a zero-shot TTS system that learns to control and model complex speech attributes, including prosody, content, and acoustic characteristics. The content attribute is modeled as illustrated in Figure 2 (Teacher box), while the DFPA generates tokens that capture expressive prosody and realistic acoustic characteristics consistent with a given speech prompt (Figure 2b, orange arrow). We then adapt this pre-trained model to the video dubbing domain, enabling it to inherit the advantages of the large-scale TTS corpus by integrating visual cues. This adaptation is achieved through two key components: (a) CCTA (Section 3.4.3), which transfers knowledge from the TTS domain to generate content representations consistent with the target text transcription, achieving precise lip synchronization and capturing global prosodic states conveyed through visual cues; and (b) DFPA modeling with modified conditioning to better suit the visual modality which is detailed in Section 3.4.4.
@@ 3.2 Data Preparation
We leverage the pre-trained FACodec [27], which consists of a Codec Encoder and a Codec Decoder. Given a speech waveform $\mathbf{S}$, we apply the Codec Encoder to decompose $\mathbf{S}$ into discrete prosody $\mathbf{x}_{p}\in\mathbb{N}^{m\times L}$, content $\mathbf{x}_{c}\in\mathbb{N}^{n\times L}$, and acoustic tokens $\mathbf{x}_{a}\in\mathbb{N}^{k\times L}$, along with a speaker embedding $\mathbf{s}\in\mathbb{R}^{D_{s}}$, where $m$, $n$, $k$ are the number of Residual Vector Quantization (RVQ) codebooks (each with vocabulary size $v$), $L$ is the token sequence length, and $D_{s}$ is the speaker embedding dimension. The Codec Decoder reconstructs the speech waveform from these tokens and the speaker embedding.
@@
Given a text transcript $\mathbf{T}$, we use a grapheme-to-phoneme converter [40] to obtain the corresponding phoneme sequence, which is mapped into a sequence of tokens $\mathbf{P}$ consisting of $N$ tokens. A Phoneme Encoder then transforms $\mathbf{P}$ into a sequence of embeddings $\mathbf{p}\in\mathbb{R}^{N\times D}$, where $D$ denotes the hidden dimension.
@@
Given a silent video $\mathbf{V}$, we follow the extraction pipeline in [11] to obtain the face-cropped video $\mathbf{V}_{\text{face}}$ and lips-cropped video $\mathbf{V}_{\text{lip}}$ and their corresponding face feature embeddings $\mathbf{v}_{\text{face}}\in\mathbb{R}^{F\times D_{f}}$ and lip feature embeddings $\mathbf{v}_{\text{lip}}\in\mathbb{R}^{F\times D_{l}}$. Here, $F$ is the number of frames, and $D_{f}$ and $D_{l}$ denote the face and lip feature dimensions, respectively. Since these extracted features exist in different representation spaces, we apply a 1D convolution to map them into a shared space: $\hat{\mathbf{v}}_{\text{face}}\in\mathbb{R}^{F\times D}$ and $\hat{\mathbf{v}}_{\text{lip}}\in\mathbb{R}^{F\times D}$.
@@ 3.3 Zero-shot TTS Pre-training
This stage aims to generate high-quality speech with natural prosody and expressive characteristics. To achieve this, we follow the zero-shot TTS setting, which aims to synthesize speech that faithfully reproduces the speech attributes of unseen speakers using only a few seconds of their reference speech. Let $\mathbf{S}^{r}$ denote the reference speech. Our objective is to build a framework that synthesizes $\tilde{\mathbf{S}}$ by generating prosody $\tilde{\mathbf{x}}_{p}$, content $\tilde{\mathbf{x}}_{c}$, and acoustic $\tilde{\mathbf{x}}_{a}$ discrete token sequences conditioned on $\mathbf{S}^{r}$ and $\mathbf{T}$. The key components of this stage are described below.
@@
Figure 2 illustrates the content attribute modeled by a simple deterministic architecture (Teacher box). The phoneme sequence $\mathbf{p}$ is processed to predict $n$ representations (one per RVQ codebook of content). These representations are then passed through a Projection to obtain a latent representation of content $\tilde{\mathbf{h}}_{c}\in\mathbb{R}^{n\times L\times D}$, and through a Content Head to produce a probability distribution over content tokens, from which $\tilde{\mathbf{x}}_{c}$ is predicted.
@@
For prosody and acoustic attributes, we design a discrete flow matching framework to model their joint distribution. As shown in Figure 2b, the Discrete Flow Denoiser outputs the joint prosody-acoustic distribution, represented as a probabilistic velocity field. Our denoiser is built on a Diffusion Transformer (DiT) [41], which takes the current denoising target $\mathbf{x}_{t}=[\mathbf{x}_{p,t};\mathbf{x}_{a,t}]\in\mathbb{N}^{(m+k)\times L}$ as input, where $[;]$ denotes channel-wise concatenation, and $\mathbf{x}_{p,t}\in\mathbb{N}^{m\times L}$ and $\mathbf{x}_{a,t}\in\mathbb{N}^{k\times L}$ are the masked prosody and acoustic token sequences at step $t\in[0,1]$, respectively. Here, $\mathbf{x}_{0}$ denotes the fully masked source sequences and $\mathbf{x}_{1}$ denotes the ground-truth token sequences. The denoiser is conditioned on $\mathbf{c}$, which comprises prosody tokens $\mathbf{x}^{r}_{p}\in\mathbb{N}^{m\times L^{\prime}}$, acoustic tokens $\mathbf{x}^{r}_{a}\in\mathbb{N}^{k\times L^{\prime}}$ (where $L^{\prime}$ is the token sequence length of $\mathbf{S}^{r}$), and a speaker embedding $\mathbf{s}$ extracted from $\mathbf{S}^{r}$ via the Codec Encoder, along with the latent representation of content $\tilde{\mathbf{h}}_{c}$. These conditioning signals provide contextual information to predict the features in $\mathbb{R}^{(m+k)\times L\times v}$ that represent the posterior distributions for prosody and acoustic tokens. This design enables the model to simultaneously generate tokens that capture expressive prosody $\tilde{\mathbf{x}}_{p}$ and realistic acoustic characteristics $\tilde{\mathbf{x}}_{a}$, remaining consistent with the speaker’s vocal identity. The objective of the DFPA module is to learn a probabilistic denoiser $p(\mathbf{x}_{1}|\mathbf{x}_{t},\mathbf{c})$ that reconstructs masked tokens under varying masking ratios. Formally, the model is trained with the Cross-Entropy loss:
@@
where $p_{1|t}(\cdot\mid\mathbf{x}_{t},\mathbf{c};\theta)$ denotes the posterior distribution of $\mathbf{x}_{1}$ given the partially corrupted sequence $\mathbf{x}_{t}$, conditioning inputs $\mathbf{c}$, and $\theta$ represents the parameters of the denoiser. Here, $\mathcal{T}=\{1,\ldots,(m+k)L\}$ represents the index set of token positions, $\mathbf{x}_{t}\sim p_{t}(\mathbf{x}\mid\mathbf{x}_{0},\mathbf{x}_{1})$ is sampled from the mixture path, and $\mathbf{x}_{0}\sim p$, $\mathbf{x}_{1}\sim q$ denote the source and target distributions, respectively.
Once trained, this TTS model serves as a teacher in our dubbing framework. Its pre-learned representations of pronunciation, prosody, and acoustic characteristics are transferred to the video-driven dubbing stage, providing strong initialization for naturalness, pronunciation accuracy, and expressive speech generation. We encourage readers to refer to the Supplementary Material for details on constructing the conditioning inputs, the model architecture, and training configurations of the zero-shot TTS pre-training stage.